OpenStudy (anonymous):

How to find the inflection points of (x^7/42)-(x^5/10)+(x^3/6)+(3x)+4? I know that it involves finding the zeros of the 2nd derivative and looking at concavity at those points, but I can't get to the zeros...

OpenStudy (anonymous):

to get POI (points of infection) find where f" is equal to 0, you should use the equation feature so i can follow the equation better

OpenStudy (anonymous):

\[((x^7)\div42) -((x^5)\div10)+((x^3)\div6)+3x+4 \]

OpenStudy (callisto):

\[f(x) = \frac{x^7}{42}-\frac{x^5}{10}+\frac{x^3}{6}+3x+4\] What is f'(x) ?

OpenStudy (anonymous):

\[x^6/7 -x^4/2 +x^2/2+3\]

OpenStudy (anonymous):

And then f double prim is \[x^5-2x^3+x\]

OpenStudy (callisto):

f'(x) is not right, f''(x) is right.

OpenStudy (anonymous):

*prime

OpenStudy (callisto):

So, we need to solve \[f''(x)=0\]\[x^5-2x^3+x =0\]Factorize the left side, what do you get?

OpenStudy (anonymous):

x(\[x^4-2x^2+1\]) ?

OpenStudy (callisto):

Factorize \(x^4 -2x^2 +1\) too..

OpenStudy (callisto):

\[(a-b)^2 = a^2-2ab+b^2\]Take a = x^2 and b = 1

OpenStudy (anonymous):

OOOOOOOOOOOOOOOOOOOH RIIIIIIIIIIIIIIIIGGGGGGGGGGGGHHHHHHHHHTT So it would be x^2 -1 ?

OpenStudy (callisto):

No. Instead, it should be \((x^2 -1)^2\). Though, you can still factorize \(x^2-1\)

OpenStudy (anonymous):

So, would the final factorized version be x((x-1)(x+1))^2 ?

OpenStudy (callisto):

Yup! :)

OpenStudy (anonymous):

THANK YOU!

OpenStudy (callisto):

Welcome :)